# TongUI: Building Generalized GUI Agents by Learning from Multimodal Web Tutorials

### v2

@@ 5.2.2 Offline Navigation
Table 7: Results on Baidu Experience.
5.2.3 Online Navigation
We evaluate the online navigation performance on the MiniWob benchmark, and the results are shown in Tab. 8. Compared with offline navigation, online navigation tasks are more challenging, since they require real-time processing of dynamic environments, handling unexpected obstacles, and adapting to changes in the navigation path. In this case, TongGUI that learns from multimodal tutorials achieves improvements again. Considering that the multimodal tutorials are offline data, this improvements confirm the generalization of TongUI.
Model FT Score
CC-Net (SL) [11] $\checkmark$ 23.4
Pix2Act [33] $\checkmark$ 55.2
SeeClick-9.6B [4] $\checkmark$ 67.0
Qwen2-VL-2B [36] $\checkmark$ 66.8
ShowUI-2B [18] $\checkmark$ 71.5
Qwen2.5-VL-3B $\times$ 0.3
Qwen2.5-VL-3B-ShowUI $\checkmark$ 67.7
TongUI-3B $\checkmark$ 72.7
TongUI-7B $\checkmark$ 73.9
Table 8: Results on MiniWob.
5.3 User Study for Data Quality
To validate the effectiveness of our data filtering strategy and evaluate the data quality of the collected GUI-Net-1M dataset, we conduct two user studies. In both studies, five participants with substantial experience in GUI agent research (but not involved in our project) are asked to rate the quality of each data point on a scale from 0 (very poor) to 5 (excellent). As shown in Fig. 6(a), we evaluate the impact of our Trajectory Filtering step. Specifically, we randomly sample 100 data points before filtering and 100 after filtering, and mix them. As shown in Fig. 6(a), the average score increases from 3.22 to 4.12 after filtering, demonstrating that Trajectory Filtering significantly improves data quality by effectively removing low-quality or GUI-irrelevant steps. We also randomly sample 100 data points from the ShowUI dataset for comparison. The results are shown in Fig. 6(b), the ShowUI dataset received an average score of 4.26. This suggests that data quality in GUI-Net-1M is comparable to that in the ShowUI dataset, while our data leads to further improvements.
Figure 6: Average scores from humans
6 Conclusion
In this paper, we present the TongUI framework that allows GUI agents to learn from multimodal web tutorials. By defining suitable and rich sources, we can crawl multimodal tutorials about diverse applications in different operating systems. The proposed tutorial processing method can extract tasks and trajectories from tutorials, and we obtain the GUI-Net-1M dataset. Using this dataset, the tuned Qwen2.5-VL model achieves improvements on multiple commonly used benchmarks, demonstrating the effectiveness of the TongUI framework and the collected GUI-Net-1M dataset. In this method, we have to collect all tutotrials and train the model once, which may overlook the potential for continual learning. In the future, we will explore the continual learning capability of GUI agents, enabling the agents to better adapt to new environments.

### v3

@@ Appendix C More Experiments
TongUI-32B 88.2 56.9
Table 16: Results on the Baidu Experience
Data Scaling Effect
We evaluate the scaling law of GUI agents using collected data from different sources with Qwen2.5-VL-7B. Concretely, (1) we report the performance of w/o SFT; (2) we use the refined public data to tune the model; (2) we add data collected from WikiHow to tune the model; (4) we add data collected from Baidu Experience to tune the model; (5) we add data collected from Video tutorials to tune the model. Results is shown in table 12. We find that with the increase of data numbers and data sources, the performance of GUI agents is also constantly improving.
Error Bar
To evaluate the robustness and stability of our model, we conduct three independent runs on Mind2Web for each type of tasks: Cross-Task, Cross-Web, and Cross-Domain, using TongUI-3B. Table 13 reports the performance on three evaluation metrics(Element Accuracy, Operation F1, and Step Success Rate) along with their mean and standard deviation (Mean ± Std) across runs. The results show that the model exhibits consistent performance with low variance, indicating stable performance of our model.
Results on AndroidControl
In table 14, we show the model performance of our model fine-tuned in the AndroidControl training set. The result indicates that our model achieves the best performance on both low and high level instructions compared to other models trained on the same training set.
Results on GUI Odyssey
In table 15, we show the results evaluated on GUI Odyssey Test-Random split. TongUI-7B achieves the best performance compared with open-source and closed-source models.
Appendix D Examples
GUI-Net
We visualize several examples of collected trajectory data, as shown in figs. 9, 10, 11, 12 and 13.
Figure 9: Dataset example.
Figure 10: Dataset example.
Figure 11: Dataset example.
Figure 12: Dataset example.
Figure 13: Dataset example.
TongUI
We visualize several examples executed by the TongUI agent, as shown in figs. 14, 15, 16 and 17.
Figure 14: Examples executed by TongUI.
Figure 15: Examples executed by TongUI.
Figure 16: Examples executed by TongUI.
Figure 17: Examples executed by TongUI.